Prev Next

SW&IT Meeting Minutes 2023-02-06

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20230123
LR20230123 - Bec, Matthieu 80.0
LR20230123 - Mechtley, Brandon 80.0
LR20230123 - Eychaner, Glenn 80.0
LR20230123 - Edmeades, David 80.0
LR20230123 - Grenz, Paul 36.0
LR20230123 - Hooper, Stephen 0.0
LR20230123 - Zhang, Yang 80.0
LR20230123 - Zhang, Xianyu 104.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Maybe IT should have a way to run a single command on every registered (Linux) computer and VM, for checking things
  • Secure System Administration and Developer Certification" enrollment available.

Team reports

MB

  • Last week
    • OC, Management, AOPM, AO.SYS.SW
    • tcs-simulator full debug with vs-code front-end. check gdb 'rr', python plugins etc.
    • tcs wlsp rpm fix
    • filter-out ack events from ahs-slack
  • This week
    • tcs-simulator, IT7830
    • telemetry timestamps, IT8671
    • ao.sys IssueTrak and github issues update
    • web-iif framework

XZ

  • Last week
    • DX ASM/AO supprt, data analysis
    • ASM SW patch release
    • ASM system and upgrade document
    • DX ASM taken down preparation
  • This week
    • ASM system and upgrade document
    • ASM/AO support
    • DX ASM issue analysis

YZ

  • Previous:
    • TCS updated (OSS/PCS/MCS config/event.dat/ARGOS light proxy).
    • Additional tests for IT8747 (MODS GUI problems) with Olga. It appears to be working fine after deleting some dot files on inaf@obs1.
    • Tweaked some MODS alert services configuration.
    • Worked on archive/LBC/MODS/TMS utilities.
  • Next:
    • Additional TCS updates?
    • Continue to work on archive/LBC/MODS/TMS utilities.

DE

  • Last week:
    • Office trip
      • ARGOS laptops, nothing from JCG about plan.
      • LUCI-new is unhappy, something to do with the RAID controller. I have troubleshooting to do tomorrow.
      • My workstation scratch drive is probably dead, lost some images....
    • Nothing so far from JCG re: PMAC laptop for me to reimage it
    • Volume6 on mountain getting awfully full; I couldn't find any obviously badly stored data
    • Alessandro looking for ex-JW laptop PW; I was leaving this for MarkS given previous interactions, especially since I don't have the JW user account PW.
    • Offered to help MarkS with Vault; my records indicate that I did this for him 2 years ago but who knows what's going on. Nothing back from him on this.
    • tcs-testvirtualization ready to test
  • This week:
    • More office for luci-new
    • Veeam maintenance renewal?
    • I put the drive from my workstation in the freezer; we'll see if the old magic still has power
    • Hopefully some feedback from JCG and MarkS

GCE

  • Last Week
    • Nearly completed scripts for installation of the LUCI and GEIRS software on luci-new.
      • Unfortunately, luci-new has had a major hardware issue.
  • This Week
    • Finish testing the LUCI and GEIRS software install scripts for luci-new.
    • Possible work on AO TCS changes (phase 2.1 and 2.2).
    • Further improvements to the LUCI AHS emailer, including an interface in the LUCI LMC to use caput() to turn off and on the various sensors.
    • Look into how best to keep luci.luci and lucix.luci databases synchronized. This is a challenge.
    • Observer GUI:
      • Jenny Power reported an issue with the observer GUI failing to update which needs investigation.
      • Pauses of active items don’t propagate into the Scheduler and have been disabled for now.
      • “More Details” and “Edit Item” menus were never finished and do not work, and have been disabled for now.

BM

  • Last week
    • NCPA
    • AO SW checkout, new release procedure (for now, will merge patches and coordinate w/aog to rename checkout to stable)
  • This week
    • test plan for rerotator presetao interruptions (IT 8762)
    • vm test lbti ncpa, daytime test plan
    • make a document for separating ao-supervisor wfs/asm/lib
    • misc. bugs (daytime conf AOS message, ccd viewer overlay msg, wfsarb memory leak) -- will inventory adsec vs wfs issues for AOSW

PSG

  • Last week
    • AZCAM testing.
    • SHARK-NIR/LBTI preset - not updating OBJNAME header info properly.
    • White light source testing - proxy works fine, OSS not updated yet.
    • Vacation day Friday
  • This week
    • IT8769 - PEPSI preset fails due to unresponsive AzCam: related to AZCAM work.
    • Rework GCS refmem
    • Add features to PMC throttling.
    • Continue fixing LBTI preset.
Topic revision: r13 - 06 Feb 2023, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback